OpenStudy (eric_d):

Expand the following expression as a series in ascending powers of x up to the terms in x^3. State the range of values of x in each case for which expansion is valid.

OpenStudy (eric_d):

@ganeshie8

OpenStudy (eric_d):



OpenStudy (eric_d):

@phi

OpenStudy (phi):

is this calculus? and do they expect a taylor series expansion?

OpenStudy (eric_d):

Binomial expansion

OpenStudy (phi):

are you sure? The binomial expansion tells us how to expand forms (a+b)^n

OpenStudy (eric_d):

Sorry, it's binomial series

OpenStudy (phi):

oh. According to https://en.wikipedia.org/wiki/Binomial_series that means taylor series about x=0 of a function in the form \[ f(x) = (1+x)^\alpha \] so the first step is write your expression in that form

OpenStudy (eric_d):

The person who taught me.. had left.. He said I can use this http://prntscr.com/49rlmz

OpenStudy (phi):

yes, but we need to write you expression in that form. Let me think about it.

OpenStudy (eric_d):

ok

OpenStudy (phi):

Here is the idea. write your problem as \[ (x+2) ( 1 - 3x)^{-\frac{1}{2}} \] now expand only (1-3x)^ (-½) up to x^3 using https://en.wikipedia.org/wiki/Binomial_series then multiply that by (x+2) , collect terms, and keep only up to x^3 terms

OpenStudy (phi):

can you expand (1-3x)^ -½ ?

OpenStudy (eric_d):

I'm stuck at this part

OpenStudy (phi):

Here is the formula what is "alpha" for your problem? what is "x" for your problem?

OpenStudy (phi):

in other words match up \[ (1+x)^\alpha \] and \[ (1+ - 3x ) ^ {- \frac{1}{2}} \] do you see alpha is -½ ? and that x in the formula matches (-3x) ?